OpenStudy (anonymous):

Can some on please help me?? Create One function, g(x), with two real irrational solutions.

OpenStudy (amistre64):

what are 2 of your favorite irrational numbers?

OpenStudy (anonymous):

I'm not sure.... This is what my teacher said to me -->I see your thinking for g(x) and h(x). They're in factored form, so setting them equal to zero and solving provides you with x = sqrt(2) and sqrt(3) for g(x) and x = +/- 2i for h(x). However, I asked that you write your answers in standard form, because g(x) is very ugly. On another note, if sqrt(2) is a solution, then -sqrt(2) is also a solution, because square roots have conjugates similar to the complex solutions. This goes for sqrt(3) as well. So your g(x) equation technically has 4 solutions, which makes it a 4th degree polynomial. Use the discriminant to make up quadratic equations with the type of solution that you want to have. The discriminant is what goes under the radical in the Quadratic Formula, so if it's a negative number, the solution is imaginary (or complex). If the discriminant is a perfect square, the solution will look like a normal number (these are rational, and include fractions). If the discriminant is not a perfect square, the answer will involve a n on-repeating decimal, which is irrational. Hopefully, this helps you rethink your answers. My first function that she told me to change is \[g(x)=(x-\sqrt{3})(x-\sqrt{2})\]

OpenStudy (anonymous):

@mathman806

OpenStudy (anonymous):

@mtmb11

OpenStudy (anonymous):

\[g(x)=(x-\sqrt{3})(x+\sqrt{3})\]